Systems Staff Meeting Notes 5/10/99

Systems Staff Meeting Notes
May 10, 1999

Present: Christian Conrad, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Everett Phillips, Suzy Shaw, Ying Tang
Announcements
  1. Christian Conrad, the new Web Administrator, will have phone extension 149. Suzy's extension will change to 150 when she moves upstairs.
Review Minutes of 5/3/99
  1. Network #1 - NERDC has been alerted - whenever they want to stop the VLAN filtering they can.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. 40 Cows (80 boxes) were received on Friday. They are currently housed in our vaults.
  2. The space in LAC is ready for the storage of old equipment. The key situation is still undecided. Richard Phillips will also need access to the room so he can check for leaks when it rains. We will probably issue a separate key for the door. There is a hang-up with the elevators, access to the first floor is the issue. This would really only present a problem after regular business hours. Just don't let the elevator doors close after 5:00 PM or you'll have to take the stairs down. We will encourage staff to return items during regular hours. They will have to send us a list of the items they are going to return, then they can load up the room on their own.
  3. Need to remind John Ashcraft to return the old machines in return for the new ones he has already picked up. These new Cows are meant as replacements for old 75 & 100's.
  4. More HD's are on order. We have none on hand at the moment. These are 10 GIG (getting bigger for less money).
  5. Network switch order was finalized at 45. These should be replacing all ports & will need IP addresses. Not all the IP's have been cleared out in order to become available for the switches - Bobbie has notified Rich that we need some Circulation access defined.
  6. When the final cost of the network wiring has been received from Digital, we will order 40 more Cows.
Networking
  1. NERDC - currently experiencing a network outage. Reported as a problem with the Criser Catalyst. The supervisor module is bad. They have moved to Fast Ether and now we are unable to get off campus. Of course, this kind of network trouble and CORE problems have been occurring this week when many NERDC staff are at the Interop Conference in Las Vegas.
  2. Network Advisory Committee is scheduled for today, will cover the current outage, state of the network & user authentication. Changes are expected.
  3. Official announcement came down last week, Earl Robbins has been made the interim CIO for UF. The position is interim for one year, at which time the need for the position will be evaluated.
  4. Wiring Project Status - MSL - outage on Friday for a short time. Techs were fumbling until they got it right.
Software Projects Status
  1. LASSIE - old information has been entered in the database. Working on the baseline database, setting up Carolyn Allen & Cynthia Bowker.
  2. A database has been set up for patron tracking in Special Collections. They are tracking whom checks out what, where & when. Special Collections would like to gain more historical information on checkouts. The upcoming user authentication process may intersect with this project.
  3. Winston & Ying have renamed the IR database to CITS. This won't do, it is already in use on campus. They will try again.
DP Coordinator
  1. Our new student, Brian, started today and was able to experience the full force of the phone features, due to the rash of phone calls reporting networking problems.
  2. Student schedules should be finalized this week, as soon as Leo shows up and turns in his work permit.
  3. Compustat tapes - Yet another version of the yearly tapes are on the way. Andy Halula notes that changes were made on their end in the Fall of '98. Bill is trying to get a list of other universities that can actually read the yearly tapes they've received. If we are unable to read this next batch, we will try swapping tapes with one of the schools on the list.
Calendar
  1. Ashley is out all this week.
  2. Brian Burton out today.
  3. Rich returns tomorrow.
  4. Brian out May 13 & 14.
  5. Rich out May 14.
  6. Brian out May 28-June 1.
  7. Bobbie & Winston are both out July 12-July 16.


Back to the Systems Homepage

Last updated February 18, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu